AnakKrakatau1883:

Simplify three square root of five end root minus two square root of seven end root plus square root of forty five end root minus square root of twenty eight. two square root of twelve two square root of two six square root of five end root minus four square root of seven six square root of ten end root minus four square root of fourteen

Tranquility:

\(\ 3\sqrt{5} - 2\sqrt{7} + \sqrt{45} - \sqrt{28}\) Does this look like your question?

Tranquility:

Here's a rule that you need to know: \(\sqrt{a^2 \times b} = a\sqrt{b}\) Knowing that, how would you simplify \(\sqrt{45}\) which is equal to \(\sqrt{9\times 5}\) which is the same as \(\sqrt{3^2 \times 5}\) How would you simplify that?

Tranquility:

Similarly, 28 is equal to 4 * 7. And you should know that 4 is 2^2. Therefore, how would you simplify \(\sqrt{28}\) ?